Musgrave Corp. has fixed costs of $46,000 and variable costs that are 30% of the current...

  1. Musgrave Corp. has fixed costs of $46,000 and variable costs that are 30% of the current sales price of $2.15 per unit. At this price, they sell 40,000 units. Musgrave can increase sales by 10,000 units by cutting its unit price from $2.15 to $1.95, but the rate of variable cost per unit will still be 30% of the new sales price.

  1. What is the change in EBIT if Musgrave decides to cut its price? (6 points)
  1. What is the DOL at BOTH sales levels?
Answer #1

a. Change in EBIT = (22,250-14,200) = Increased by 8,050

Particulars Selling price $ 2.15 Selling price $ 1.95
Selling price per unit (a)                           2.15                           1.95
Number of units (b)                  40,000.00                  50,000.00
Total Sales (c=a*b)                  86,000.00                  97,500.00
Varibale cost (d=c*30%)                  25,800.00                  29,250.00
Fixed Cost (e)                  46,000.00                  46,000.00
EBIT (c-d-e)                  14,200.00                  22,250.00

*variable cost also reduced to 30% of new price

Degree of Operating Leverage = (% change in EBIT / % Change in Sales)

= (56.69%/13.37%)

= 4.24

Total Sales @ 2.15 (a)                  86,000.00
Total Sales @ 1.95 (b)                  97,500.00
% Change in Sales (((b-a)/a)*100) 13.37%
EBIT @ 2.15 (c)                  14,200.00
EBIT @ 1.95 (d)                  22,250.00
% Change in EBIT (((d-c)/c)*100) 56.69%